Use the aggregate expenditures model and the following values to answer the next question. MPC $900 0.9 $2,500 $2,500 $1,000 Determine equilibrium consumption for this economy. Multiple Choice $44.100 $45,000 $45,900 $50,000
Answer #1

Answer : Option B:$45000

Let Y is equilibrium income level

Y=C+I+G

C=A+MPC*(Y-T) =900+0.9*(Y-1000)=0.9*Y

Y=0.9Y+2500+2500

Y=$50000

So equilibrium consumption C=0.9*Y=0.9*50000=$45000